LEP EXPERIMENTS COMMITTEE
Minutes of the 47th meeting,
11 November 1997

OPEN SESSION

1. Future LEP operation

L. Foà reported on the current situation regarding operation of LEP in 2000. This is strongly supported by the Research Board and by the SPC, which defines it to be the top near-term scientific priority for the laboratory. However, in view of the severe financial constraints, special contributions will be necessary. These have been requested from the Member and non-Member States and, so far, there has been an encouraging yet incomplete response. After the remaining funds are secured, the final decision is expected in March 1998.

2. LEP2 status

R. Bailey reported on the status of LEP2. Despite the 35 day delay in LEP startup after the fire in the SPS surface building, 1997 has been a record year for LEP performance. A total of 73.3 pb-1 per experiment has been delivered, comprising
63.8 pb-1 at beam energies of 90.5-92 GeV, 7.2 pb-1 at 65-68 GeV and 2.3 pb-1 at 45 GeV. Operating with 90/60 optics and 4 bunches per beam, LEP reached a maximum current of 5.2 mA, a peak luminosity of 5x1031 cm-2s-1 and a maximum integrated luminosity of 1.8 pb-1 per day. LEP achieved a good average betatron coupling parameter (emittance ratio), k=0.008, and a large maximum beam-beam tune shift of 0.056. Furthermore a beam polarization of 5-12% was measured at beam energies between 41.2 and 55.3 GeV, with 1-2% polarization observed at 60.6 GeV. This is the highest energy so far for polarized beams, and is an encouraging sign for LEP2 beam energy calibration.

At the end of the cycle a successful test was made of the new 102/90 low emittance optics and the results appear promising for high energy operation. The machine was rapidly re-commissioned after re-cabling and achieved a peak luminosity of 4.5x1031 cm-2s-1 at a record beam energy of 92 GeV. Detector backgrounds were low.

The increased energy and high currents have caused some problems. Some detectors have experienced high backgrounds - notably DELPHI during injection and OPAL during physics. The mean time between trips of a sc rf module is 1 h (but less than 10% of fills are actually lost due to rf trips). Several vacuum chamber leaks have been experienced due to synchrotron radiation heating. Many of the transition regions where this heating occurs have been smoothed, and further improvements are presently underway, e.g. removal of the pretzel separators.

During the 97-98 shutdown another 8 sc rf modules (32 cavities) will be installed and more of the warm copper cavities will be removed. This will allow the operating energy in 1998 to be raised near to ÷s = 189 GeV and, based on this year's experience, an average integrated luminosity in excess of 1 pb-1 per day is realistic. With 117 days expected for LEP physics at high energy in 1998, the total luminosity could exceed 120 pb-1. Better estimates of luminosity and beam energy await the outcome of Chamonix workshop in January.
 

3. Reports on the LEP experiments

               L3:          M. Pohl
              OPAL:     A. Honma
              ALEPH:   P. Dornan
              DELPHI:  P. Charpentier

Each experiment reported samples of about 700 WW events in 1997, which provide a statistical precision on the W mass of about 180 MeV per experiment. When the four experiments are combined, the resulting precision of 90 MeV matches the present measurement from the Tevatron.

Cross section upper limits of 0.2-0.5 pb have been obtained for the production of SUSY particles. Although the new data have revealed no clear evidence for new physics, numerous "interesting" candidate events were presented - especially in the Higgs sector. At ÷s = 183 GeV, LEP is sensitive to a Higgs mass up to about 90 GeV in the Higgs-strahlung channel, HZ, when the individual limits of 82-84 GeV (95% C.L.) are combined. However LEP has just crossed the threshold for ZZ pair production, which generates significant backgrounds in this mass region. Higgs and Z boson events can be distinguished from each other by the more abundant b quarks in Higgs decays. The experiments showed clear examples of ll-bb and jj-bb events which, although compatible with ZZ events, match the characteristics of HZ events. It is clear that much more statistics are needed to explore this interesting mass range.

The detectors reported some operational problems. L3 has a short in one of the forward toroid coils, which is presently under repair. OPAL reported on a wire breakage in the stereo cells of the vertex chamber which has resulted in a 33% loss; no repair is being made, however, since this has a negligible overall effect on the data. DELPHI reported a drop in efficiency of the forward muon chambers from 95% to 70%, for which the cause is under investigation. On a different note, ALEPH announced a software tracking upgrade which provides efficiency gains of 5-40%, improved mass resolutions (up to 8%) and improved vertex resolution.

Concerning the future operation of LEP, the four collaborations unanimously repeated their enthusiasm and strong support for operation of LEP in 2000 at an energy near ÷s = 200 GeV.

4. LEP working group reports

             Four-jet events:   D. Schlatter
             Supersymmetry:   L. Pape

D. Schlatter reported on the outcome of the special run of 6 pb-1 per experiment at 130-136 GeV to resolve the 4-jet puzzle. No excess in the 105 GeV mass region has been observed in the new data by any experiment, which excludes a signal of 1.5 pb at 99% C.L. (i.e. one sigma on the conservative side of the previously-measured cross section, 2.2±0.7 pb). The conclusion is that new physics is ruled out as an explanation of the ALEPH 4-jet events.

L. Pape reported on the preliminary results of the Supersymmetry working group. From the combined LEP data, cross section upper limits of about 0.1-0.2 pb for the production of SUSY particles are obtained, as well as improvements relative to individual experiments of 5-10 GeV in the lower mass limits.

CLOSED SESSION

Present: S. Bethke, J. Drees, L. Foà, F. Gasparini, K. Hübner, G. Kantardjian, J. Kirkby (Secretary),
                 K.-H. Kissler, W. Lohmann, T. Lohse, M. Mangano, R. Marshall, T. Müller, S. Myers,
                 L. Pape, S. Pokorski, Y. Sirois, P. Sphicas, I. Videau, P. Wells and P. Zerwas (Chairman).

Apologies: J. Dainton, G. Goggi, P.O. Hulth, J. May, J. Panman, M. Turala.

1. Approval of the minutes of the 46th meeting

The minutes of the 46th meeting (LEPC 97-9/LEPC 46) were approved without modification.

2. Chairman's report

The Chairman and Research Director welcomed P. Sphicas who has accepted to take over as LEP Coordinator from P. Wells, starting 1 April 1998. The Chairman announced that D. Stickland has agreed to chair the new LEP Archive working group which comprises two representatives from IT Division and from each of the four LEP experiments.

3. Discussion on the LEP machine reports

The Chairman joined with all the members of the LEPC in congratulating the machine team for the record performance of LEP in 1997. The new 102/90 optics is especially encouraging since it not only allows good performance at high energy but also efficient and compatible operation at lower energies near the Z (using 90/90 optics), without any re-cabling. The committee looks forward to hearing the performance projections for 1998 at its next meeting, following the Chamonix workshop in January.

4. Discussion on the reports from the LEP experiments and working groups

The committee discussed the Higgs-strahlung channel, HZ, which must be disentangled from the significant background channel, ZZ. Since b quarks are readily produced in Z decays, large data samples are needed to search for Higgs bosons near the Z mass. Of particular interest is the lightest SUSY Higgs boson, which could have a mass in this region and for which the new data are eagerly awaited.

The committee was pleased to see a resolution of the 4-jet puzzle from the dedicated one-week run at 130-136 GeV. Although the central question of new physics has now been ruled out as the source of these events, no explanation other than a statistical fluctuation has been identified.

The committee noted with interest the improvements in the mass limits of supersymmetric particles from the combined LEP experiments. A brief discussion took place on the guidelines for combining data by the LEP working groups. Now that results are starting to appear, it was felt that the successful experience of the long-established electroweak working group should be transmitted to the new groups, in order to ensure a common approach which is agreed by all the collaborations. A small group which includes the working group chairmen will be asked to formulate a brief set of guidelines, in consultation with the LEP experiments.

5. Report from the LEP Coordinator

The LEP Coordinator reported that LEP operated for 114 days in 1997, representing a loss of only 18 days relative to the original plan prior the SPS building fire. The average luminosity over the entire period at high energy has been 0.9 pb-1 per day, with a consistently higher average during the latter part of the year.

In 1998 LEP is expected to operate for 157 days in the period from 11 May until 1 November, comprising about 10 days at the start of the cycle for 2.5 pb-1 of Z calibration data, about 117 days for high energy running (÷s = 189 GeV) and about 30 days for machine development.

Concerning the beam energy calibration at LEP2, the flux loop/NMR measurements of the LEP dipoles have now been calibrated at four energies between 41 and 55 GeV by resonant depolarisation measurements. Averaging all measurements, and assuming a linear extrapolation, these indicate an error in the beam energy of 10±5 (stat) MeV
at 90 GeV. However individual measurements indicate different slopes, and more studies are necessary before firm estimates of errors can be made.

The LEP Energy working group is also investigating independent methods that can directly measure the beam energy at high energy. These include radiative returns to the Z (which, however, appear to have insufficient precision), Moller scattering and a precision spectrometer. Concerning the spectrometer, a dedicated dipole and associated precision beam position monitors are planned for installation at LEP during the present shutdown.

The committee expressed its support for the LEP Energy working group to find a suitable method for making a direct measurement of the beam energy at high energy. There is a clear interest to have an independent method that can be used to cross-check and improve the precision of the present technique based on extrapolation of the flux loop/NMR measurements. The committee looks forward to hearing further progress at its next meeting.

6. Any other business

The Chairman announced that the LEP jamboree of results from the 1997 run will be combined with the next LEPC meeting on 31 March.

7. Next LEPC meeting

The dates of the next meeting of the LEPC are Tuesday-Wednesday, 31 March - 1 April. The tentative dates for the LEPC in the remainder of 1998 are as follows:

               15 - 16 September
               12 - 13 November

                                                                                                                 J. Kirkby

DOCUMENTS RECEIVED

[1] LEP Higgs Working Group, Lower bound for the SM Higgs boson mass: combined result from the four LEP experiments, LEPC 97-11.